OpenStudy (anonymous):

A wooden pyramid, 9 inches tall, has a square base. A carpenter increases the dimensions of the wooden pyramid by a factor of 4 and makes a larger pyramid with the new dimensions. Describe in complete sentences the ratio of the volumes of the two pyramids.

OpenStudy (anonymous):

@waterineyes

OpenStudy (anonymous):

See here you are given with: Height of Pyramid that is 9 in. Now the formula for Volume of Pyramid is: \[Volume = \frac{1}{3} \times l \times w \times h\]

OpenStudy (anonymous):

okay, so what do I do?

OpenStudy (anonymous):

And one more thing, square base means length is equal to width of that pyramid.. So, l= w. \[Volume = \frac{1}{3} \times 9 \times l^2\]

OpenStudy (anonymous):

cuz I don't have the rest of the equation

OpenStudy (anonymous):

so I just solve for that?

OpenStudy (anonymous):

Can you simplify the above that I gave you?? Keep l^2 as it is there, just divide 9 by 3 there, what you got?

OpenStudy (anonymous):

volume = 9 * l^2

OpenStudy (anonymous):

9/3 = ??

OpenStudy (anonymous):

oh crap sorry im meant to put 3

OpenStudy (anonymous):

volume = 3 * l^2

OpenStudy (anonymous):

Yep.. So this is the original volume, let us say it as 1, means at base of volume and l, just put 1 like this: \[\large Volume_1 = 3 \times l^2_1\] Getting?

OpenStudy (anonymous):

Idk :0

OpenStudy (anonymous):

Just put 1 in the base to show that this is the volume of first pyramid, because now pyramid has been modified by changing its base..

OpenStudy (anonymous):

so... volume = 3 * 1???

OpenStudy (anonymous):

In simple language, if you have two pyramids, then how will you differentiate their volumes?? You will let \(V_1\) as volume of first pyramid and you will suppose \(V_2\) as volume of second pyramid. Right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Hey, \(Volume_1\) and \(l_1\) is just showing that these are for first pyramid. Similarly now we will find for second case..

OpenStudy (anonymous):

Now, in question the square base has been changed to 4 times . Means l, w and h are increased by factor of 4.. So: \[l_2 = 4l_!\] \[w_2 = 4w_1\] \[h_2 = 4 h_1\]

OpenStudy (anonymous):

okay yu are confusing me

OpenStudy (anonymous):

ALOT

OpenStudy (anonymous):

What I think, you always confuse between why we take 1 and 2 in bases..

OpenStudy (anonymous):

es I don't understand the sub 1 and 2

OpenStudy (anonymous):

*yes

OpenStudy (anonymous):

Hey they are just Volume, length, width and height but sub in 1 shows that they are for first pyramid.. Now what happen when I will changed their length, width and height, automatically volume will change now.. So I let Volume, length, width , height again but with different base ie 2 because now the previous pyramid has changed and we have one newer and second pyramid..

OpenStudy (anonymous):

And one thing you will not agree to me, but I don't know how to draw pyramid, so I can't make you understand practically by doing drawings here..

OpenStudy (anonymous):

okay so the 1 and 2 represent pyramid 1 and pyramid 2. right?

OpenStudy (anonymous):

Yep... Similarly we have two pyramid in our question.. Now can I let \(V_1, l_1, w_1, h_1\) for first pyramid?

OpenStudy (anonymous):

okay I understand

OpenStudy (anonymous):

So, In first pyramid you are given with: Volume = \(V_1\) \(l_1 = w_1\) \(h_1 = 9\) Getting this much?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

So just apply the formula here for Volume: \[Volume = \frac{1}{3} \times length \times width \times height\] \[V_1 = \frac{1}{3} \times l_1 \times w_1 \times h_1\] \[V_1 = \frac{1}{3} \times l_1 \times w_1 \times 9\] Getting?

OpenStudy (anonymous):

Here \(w_1 = l_1\), so can I write \(w_1\) as \(l_1\)??

OpenStudy (anonymous):

\[V _{1} = 3 \times l _{1} \times w\]

OpenStudy (anonymous):

* \[w _{1}\]

OpenStudy (anonymous):

But \(w_1 = l_1\) Right?

OpenStudy (anonymous):

is that right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Yep that is right but can be changed further..

OpenStudy (anonymous):

so can you just put\[l _{1}^{2}\]

OpenStudy (anonymous):

So tell me: \[V_1 =3 \times l_1 \times l_1\] Right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Yep.. So you got: \[V_1 = 3 l^2_1\] Remember this.. Now we will move to the second pyramid.. Can I let \(V_2, l_2, w_2, h_2\) for second pyramid now?

OpenStudy (anonymous):

okay go it

OpenStudy (anonymous):

*got

OpenStudy (anonymous):

So: Now you are given with that the length, height and width of the earlier that is first pyramid are changed and they are each multiplied by 4 to make new pyramid.. Getting?

OpenStudy (anonymous):

Meaning \(l_1, h_1, w_1\) all are multiplied by 4 to get new pyramid.

OpenStudy (anonymous):

\[V _{2} =12 \times l _{2}^{6} ??\]

OpenStudy (anonymous):

Good try but no..

OpenStudy (anonymous):

:(

OpenStudy (anonymous):

I mean now: \[l_2 = 4 l_1\] \(w_2 = 4w_1\) \(h_2 = 4h_1\) Use this to find \(V_2\).. :)

OpenStudy (anonymous):

ohmygosh. im on a timed practice test and my time is almost up!

OpenStudy (anonymous):

Sorry I can't do anything in it..

OpenStudy (anonymous):

\[V _{2} = 4l _{1} \times 4w _{1} \times 4h _{1} ?????\]

OpenStudy (anonymous):

You are missing 1/3 there..

OpenStudy (anonymous):

ok I give up. thx for trying to help

OpenStudy (anonymous):

You need to be more strong. Don't ever give up.. You were just near the answer..

OpenStudy (anonymous):

Thanks for tolerating me this long.. :)

OpenStudy (anonymous):

no im not! I have been working on this one problem for over an hour nd I still don't get it. im just getting more and more confused

OpenStudy (anonymous):

haha thanks for tolerating me